3) For cache design with following characteristics, complete the following tables for other parameters. Consider 16-bit for each word.
Answer #1

For 4G:

Cache Capacity for data in Byte = 2^3= 8 bytes.

Tag field = 16 bit - (12+3) bit = 1 bit

Cache Line size=2^12 /16=2^8=256 bits

For 8G:

Cache Capacity=2^4=16 bytes

For 16G:

Tag field =14 bit

Offset field =3 bit

Cache capacity=2^3 =8 bytes

Cache line size=64bit/2*16=2 (words)

Index field=log2(2)=1 bit


answered by: ANURANJAN SARSAM
Answer #2

olution 4GB Cache line size-2offset bits 25 bytes 16words Cache size 2A32 bytes Cache line size: 16 words 216 bytes #otlines-Main MemoryTag Field Index field Offset Field Cache Line Cache Capacity Capacity 4G SG 16G 9 bit 14 bit 16 bit 18 bit 14 bit